Bean Ldt is a division of Earl enterprise and has been allocated R5m for capital expansion in the forth-coming year.The management of Bean believes that the company must spread its risk by investing in projects with different risk profiles and has identified two possible investments.The capital to Bean is sufficient to invest in only one of the projects.Economic growth 0%,3%and6%.probability of occurance 0.3,0.4and0.3.Estimated return:project1 14,10and 8 Project2:8,18and22.Existing investment 6,12and16.Book value R5m, R5m and R10m.Market value R5m,R5and R15m.The division manager has requested the accounted to determine which of the two projects should be accepted using the portfolio theory. Required:1.calculate which investment should be selected in line with the portfolio management theory. 2.identify and describe the kind of risk the management of Bean Ltd wishes to spread by investing in different investments and state whether they should be concerned about reducing such risk.
1
Expert's answer
2020-07-24T10:34:57-0400

1.)   Expected returns of the two projects will be calculated as follows under portfolio management

ERp1 = R1P1 + R2P2 + R3P3

Where R = Estimated project return

           P = Probability of occurrence

Estimated Return for Project 1 = (0.3*14) + (0.4*10) + (0.3*8) = 11%

Estimated Return for Project 2 = (0.3*8) + (0.4*18) + (0.3*22) = 16%

Therefore, the investment under project 2 should be chosen since it brings a higher and better return and its more efficient.


2.)   The kind of risk that the management of Bean Ltd wishes to spread is called unsystematic risk which is always diversifiable. This risk is distinct to a firm, industry, market or economy and its usual sources are the financial and business risk.


The management should not be concerned since reducing such risk will decrease your portfolios sensitivity to market shifts, hence limiting the effect of losses that may arise as a result. Furthermore, the idea to invest in various assets ensures that they don’t all get affected in the same way by the market changes.